40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 10 14:57:01 2019, Koji, Summary, CDS, Channel recipe for iscaux upgrade D990694-B.pdfD990543A1.PDF
    Reply  Thu Jul 11 12:42:35 2019, gautam, Summary, CDS, P2 interface board IMG_7728.JPG
       Reply  Thu Jul 11 13:08:36 2019, Chub, Summary, CDS, P2 interface board 
          Reply  Thu Jul 11 13:09:22 2019, gautam , Summary, CDS, P2 interface board 
       Reply  Sat Jul 20 11:57:39 2019, gautam, Summary, CDS, P2 interface board IMG_7773.JPG
    Reply  Thu Jul 18 00:51:52 2019, Koji, Summary, CDS, iscaux electronics modifications LSC_whitening2.jpgLSC_whitening.jpg
       Reply  Thu Jul 18 22:34:40 2019, Koji, Summary, CDS, iscaux electronics modifications 
Message ID: 14775     Entry time: Thu Jul 18 22:34:40 2019     In reply to: 14770
Author: Koji 
Type: Summary 
Category: CDS 
Subject: iscaux electronics modifications 

The whitening filter modules have been restored to the crates. The SMA cables have been restored and fastened by a spanner. The ribbon cable to the antialiasing board was also connected. The backplane cables have not been moved from the upper DIN96 connector to the lower one.

Everything is expected to be good, but just keep eyes on the LSC signals as the boards were not quantitatvely tested yet. If you find something suspicious, report on the elog.

ELOG V3.1.3-